Let S be a set, R a ring and f be a 1-1 mapping of S onto R. Define + and · on S by: )) x y f (f(x)) f(y 1 + = + − )) x y f (f(x) f(y 1 ⋅ = ⋅ − . ∀ x, y∈S Show that ) (S, +, ⋅ is a ring isomorphic to R. 


1
Expert's answer
2021-04-13T13:48:06-0400

Let "S" be a set, "R" a ring and "f" be a 1-1 mapping of "S" onto "R". Define "+" and "\u00b7" on "S" by:

"x+ y=f^{-1}(f(x)+_Rf(y))", "x\cdot y=f^{-1}(f(x)\cdot_Rf(y))" for all "x, y\in S". Let us show that "(S, +, \u22c5)" is a ring isomorphic to "R". Taking into account that

"f(x+ y)=f(f^{-1}(f(x)+_Rf(y)))=f\circ f^{-1}(f(x)+_Rf(y))=f(x)+_Rf(y)" and

"f(x\cdot y)=f(f^{-1}(f(x)\cdot _Rf(y)))=f\circ f^{-1}(f(x)\cdot _Rf(y))=f(x)\cdot _Rf(y)", we conclude that "f:S\to R" is a ring homomorphism. Since "f" is bijection, "f" is also a ring isomorphism. It follows that "(S, +, \u22c5)" is isomorphic to the ring "R", and hence "(S, +, \u22c5)" is also a ring.